OpenStudy (anonymous):

If the height of a cylinder is tripled, but the area of the base stays the same, what happens to the volume? The volume doubles. The volume triples. The volume is four times greater. The volume is nine times greater. What happens to the volume of a square pyramid if the base dimensions remain the same, but the height is cut to one third of the original? The volume becomes 1/9 of original. The volume becomes 1/3 of original. The volume becomes 3 times larger than original. The volume becomes 9 times larger than original.

OpenStudy (anonymous):

1. volume triples

OpenStudy (anonymous):

:3thanks

imqwerty (imqwerty):

2) one third original

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no its 1/9 lmao

OpenStudy (anonymous):

The volume of a square pyramid is V = (1/3)Ah where A = area of base and h = height For a pyramid with base area A and height h, V = (1/3)Ah If you now substitute h with h/3, you get V = (1/3)(Ah/3) = (1/9)Ah (1/9)Ah is one-third of (1/3)Ah

OpenStudy (anonymous):

boom logic ^^^

imqwerty (imqwerty):

The volume, V, of a square based pyramid in cubic units is given by V=1/3 x A x h where A is the area of the base and h is the height of the pyramid.

imqwerty (imqwerty):

original volume of pyramid = 1/3 x A x h nw the height is 1/3rd of original volume = 1/3 x A x h/3 =(1/3 x A x h)/3 = original volume/3

OpenStudy (anonymous):

wrong @imqwerty

imqwerty (imqwerty):

tell me where am i wrong??

OpenStudy (anonymous):

V = (1/3)(Ah/3) = (1/9)Ah

imqwerty (imqwerty):

ok we have original volume 1/3(Ah)

imqwerty (imqwerty):

now the height becomes one third so height=h/3

imqwerty (imqwerty):

new volume= 1/3 Ah x 1/3 =original volume /3